Find and graph a linear regression equation that models the data.

  Temperature (oC)

No. of Ice cream sold

26

345

27

322

28

357

29

423

30

435



1
Expert's answer
2021-11-11T12:27:31-0500




"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c}
 & X & Y & XY & X^2 & Y^2\\
 \hline
 & 26 & 345 & 8970 & 676 & 119025\\
 & 27 & 322 & 8694 & 729 & 103684\\
 & 28 & 357 & 9996 & 784 & 127449\\
 & 29 & 423 & 12267 & 841 & 178929\\
 & 30 & 435 & 13050 & 900 & 189225\\
 Sum= & 140 & 1882 & 52977 & 3930 & 718312\\
\end{array}""\bar{X}=\dfrac{1}{n}\sum_iX_i=\dfrac{140}{5}=28"

"\bar{Y}=\dfrac{1}{n}\sum_iY_i=\dfrac{1882}{5}=376.4"

"SS_{XX}=\sum_i(X_i-\bar{X})^2=10"

"SS_{YY}=\sum_i(Y_i-\bar{Y})^2=9927.2"

"SS_{XY}=\sum_i(X_i-\bar{X})(Y_i-\bar{Y})=281"

"m=slope=\dfrac{SS_{XY}}{SS_{XX}}=\dfrac{281}{10}=28.1"

"n=\bar{Y}-m\bar{X}"

"=376.4-28.1(28)"

"=-410.4"

Therefore, we find that the regression equation is:


"Y=-410.4+28.1X"



Correlation coefficient:


"r=\dfrac{SS_{XY}}{\sqrt{SS_{XX}}\sqrt{SS_{YY}}}"

"=\dfrac{281}{\sqrt{10}\sqrt{9927.2}}=0.8918523"

We have strong positive correlation.

The regression equation is:


"Y=-410.4+28.1X"